OpenStudy (anonymous):

The ordered pairs (2, -11) and (5, -20) are solutions to which of the following equations? y = 3x - 5 y = -3x + 5 y = 3x + 5 y = -3x - 5

OpenStudy (anonymous):

Do you know where to start? @Anon101

OpenStudy (anonymous):

well i know that you have to put the x in the equation

OpenStudy (anonymous):

maybe i can do it and you can check my answer?

OpenStudy (anonymous):

Yes the 2 is the X so you put 2 where the variable X is at and then solve the equation. @Anon101 and sure ok

OpenStudy (anonymous):

y = -3x - 5 y = -3x2 - 5 = -11 yes y = -3x5 - 5 = -20 yes (2, -11) (5, -20)

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

I belive its the last one because if you plug 2 into where the X is at then multiply -3 times 2 you would get -6 then subtract 5 you would get -11. Which in the coordinate pairs y does equal -11. And if you substitute for the next 2 numbers they will equal out too :) @Anon101

OpenStudy (anonymous):

Yup! :) @Anon101

OpenStudy (anonymous):

thanks! :D